# Clustering Properties of Self-Supervised Learning

### v1

@@ Abstract
Self-supervised learning (SSL) methods via joint embedding architectures have proven remarkably effective at capturing semantically rich representations with strong clustering properties, magically in the absence of label supervision. Despite this, few of them have explored leveraging these untapped properties to improve themselves. In this paper, we provide an evidence through various metrics that the encoder’s output encoding exhibits superior and more stable clustering properties compared to other components. Building on this insight, we propose a novel positive-feedback SSL method, termed Representation Soft Assignment (ReSA), which leverages the model’s clustering properties to promote learning in a self-guided manner. Extensive experiments on standard SSL benchmarks reveal that models pretrained with ReSA outperform other state-of-the-art SSL methods by a significant margin. Finally, we analyze how ReSA facilitates better clustering properties, demonstrating that it effectively enhances clustering performance at both fine-grained and coarse-grained levels, shaping representations that are inherently more structured and semantically meaningful.
@@ 1 Introduction
Self-supervised learning (SSL) has emerged as a transformative paradigm in universal representation learning (Bachman et al., 2019; He et al., 2020; Chen et al., 2020a; Bao et al., 2021; He et al., 2022; Oquab et al., 2023; Assran et al., 2023), consistently surpassing supervised learning in downstream performance. Joint embedding architectures (JEA), in particular, aim to learn invariance of the same data under different transformations and noise (Bachman et al., 2019; He et al., 2020; Chen et al., 2020a), with demonstrated exceptional effectiveness in visual representation learning. Although such a pretext task may intuitively seem unrelated to capturing semantic relationships, extensive studies (Caron et al., 2021; Assran et al., 2022b) have demonstrated the strong correlation between its learned representations and semantic information.
@@
  • How to leverage clustering properties? We propose a novel SSL method, termed Representation Soft Assignment (ReSA), which employs an online self-clustering mechanism to leverage the model’s inherent clustering properties, thereby facilitating positive-feedback learning. Standard experiments demonstrate that ReSA surpasses existing state-of-the-art methods in both performance and training efficiency.
@@ 2.1 Self-Supervised Learning
In particular, the JEA usually consist of an encoder $E_{\theta_{e}}(\cdot)$ and a projector $G_{\theta_{g}}(\cdot)$. As shown in Figure 2, their outputs $\mathbf{H}=E_{\theta_{e}}(\mathbf{X})\in\mathbb{R}^{d_{e}\times m}$ and $\mathbf{Z}=G_{\theta_{g}}(\mathbf{H})\in\mathbb{R}^{d_{g}\times m}$ are referred to as encoding and embedding, respectively (where $m$ is the mini-batch size, $d_{e}$ and $d_{g}$ are their corresponding feature dimensions). Under this notation, we have $F_{\theta}(\cdot)=G_{\theta_{g}}(E_{\theta_{e}}(\cdot))$ with learnable parameters $\theta=\{\theta_{e},\theta_{g}\}$. The projector has been demonstrated to be a significant component, which can prevent invariance optimization from degrading the useful information in encoding (Chen et al., 2020a; Gupta et al., 2022). Consequently, the encoding is usually utilized as the optimal representation for downstream tasks.
@@
The main challenge with JEA is representation collapse, where both branches produce identical and constant outputs regardless of the inputs. Numerous paradigms have been proposed to avoid collapse, including contrastive learning methods, e.g. SimCLR (Chen et al., 2020a), MoCos (He et al., 2020; Chen et al., 2020b, 2021), MEC (Liu et al., 2022) that attract different views from the same image (positive pairs) while pushing apart different images (negative pairs), and non-contrastive approaches, e.g. BYOL (Grill et al., 2020), SwAV (Caron et al., 2020), DINO (Caron et al., 2021), INTL (Weng et al., 2024) which directly align positive targets without incorporating negative pairs. Although research (Ben-Shaul et al., 2023) has demonstrated that the encodings learned through SSL are highly correlated with semantic classes and exhibit strong clustering capabilities, few methods have leveraged this clustering ability to facilitate positive-feedback learning. A closely related work (Ma et al., 2023) exploits the encoding’s augmentation robustness to re-weight the positive alignment in the SSL objective functions; however, it still overlooks the rich clustering information inherent in the encodings.
2.2 What is Optimal Representation for Clustering
Ben-Shaul et al. (2023) find that within the encoder of JEA, the clustering ability of features improves progressively as intermediate layers get deeper. However, it remains unclear whether the projector exhibits a similar trend. In fact, embedding tends to be less influenced by clustering-irrelevant features, and existing cluster-based SSL methods, e.g. DeepCluster-v2 (Caron et al., 2018), SwAV (Caron et al., 2020), and MSN (Assran et al., 2022b), impose clustering constraints on embedding to prevent collapse. These observations raise an important question: Which JEA component provides the optimal representation for clustering? Is it the encoding, embedding, or perhaps the hidden layer outputs within the projector?
@@ 3 Exploring Clustering Properties of SSL
To quantitatively evaluate the clustering performance of these components, we employ widely recognized metrics such as the Silhouette Coefficient (SC) (Rousseeuw, 1987) and Adjusted Rand Index (ARI) (Hubert & Arabie, 1985). In particular, a larger mean value of SC ($\text{SC}_{\text{mean}}$) indicates stronger local clustering capability in the representation, and a smaller standard deviation ($\text{SC}_{\text{std}}$) reflects better stability in local clustering. Meanwhile, higher ARI values correspond to enhanced global clustering properties. The detailed introduction of these metrics can be found in the Appendix A.
@@
Figure 5: The framework of Representation Soft Assignment (ReSA). Here, no grad. denotes that the operation does not involve gradient propagation, norm signifies that each sample is $L_{2}$-normalized to compute cosine similarities, and sinkhorn refers to the Sinkhorn-Knopp algorithm used for clustering assignment.
@@ 4 Leverage Clustering Properties for Positive-Feedback Learning
Based on above analyses, we design a novel positive-feedback SSL method, which derives Representation Soft Assignment (ReSA) to guide the loss function among embeddings $\mathbf{Z}$ and $\mathbf{Z}^{\prime}$. See Figure 5 for the clear framework.
4.1 Online Self-Clustering ReSA
@@
We follow SwAV which uses only 3 iterations and sets the regularization parameter $\epsilon=0.05$. This algorithm does not involve gradient propagation, enabling it to be efficiently implemented on GPUs (Caron et al., 2020). After obtaining the doubly stochastic matrix $\mathbf{A}_{\mathbf{H}}$ as the representation soft assignment, it can naturally be utilized to guide relationship between the embeddings $\mathbf{Z}$ and $\mathbf{Z}^{\prime}$. Specially, we use the cross-entropy loss to promote the learning process:
$\ell_{\text{ReSA}}=-\frac{1}{2m}\bigg{(}$ $\sum_{i,j}\mathbf{A}_{\mathbf{H}}\circ\log\mathcal{D}(\mathbf{S}_{\mathbf{Z}})+$ (2)
$\sum_{i,j}\mathbf{A}_{\mathbf{H}}^{\top}\circ\log\mathcal{D}(\mathbf{S}_{\mathbf{Z}}^{\top})\bigg{)}$
where $\mathcal{D}(\mathbf{S}_{\mathbf{Z}})=\frac{\exp(\mathbf{S}_{\mathbf{Z}}/\tau)}{\exp(\mathbf{S}_{\mathbf{Z}}/\tau)\mathbf{1}_{m}}$ and $\mathcal{D}(\mathbf{S}_{\mathbf{Z}}^{\top})=\frac{\exp(\mathbf{S}_{\mathbf{Z}}^{\top}/\tau)}{\exp(\mathbf{S}_{\mathbf{Z}}^{\top}/\tau)\mathbf{1}_{m}}$ are respectively row-wise and column-wise probability distribution derived from the cosine similarity matrix $\mathbf{S}_{\mathbf{Z}}=\mathbf{Z}^{\top}\mathbf{Z}^{\prime}$ through the softmax function. $\tau$ is a scalar temperature hyperparameter and $\circ$ stands for Hadamard product.
@@
where $\mathbf{C}\in\mathbb{R}^{d_{g}\times K}$ is the prototype matrix learned by back-propagation, and $\mathbf{Q}=\text{sinkhorn}(\mathbf{Z}^{\top}\mathbf{C})$ is the cluster assignment using Sinkhorn-Knopp algorithm. The key differences and advantages of our ReSA compared to SwAV (and DINOv2) can be summarized as follows: (1) ReSA computes clustering assignments on encoding with high-quality clustering properties, whereas SwAV performs it on the less stable embedding. (2) SwAV requires learnable prototypes, which often necessitate complex design strategies, such as freezing prototypes during the early stages of training and use a large number of prototypes $K$ to ensure stability, whereas ReSA directly extracts clustering information from the representations. (3) SwAV executes the Sinkhorn-Knopp algorithm multiple times, corresponding to the number of augmented views. In contrast, ReSA only requires a single execution of this algorithm regardless of the number of augmented views. This highlights the efficiency of ReSA, particularly under multi-crop scenarios.
@@
$\ell_{\text{InfoNCE}}=-\frac{1}{2m}\bigg{(}\sum_{i,i}\log\mathcal{D}(\mathbf{S}_{\mathbf{Z}})+\sum_{i,i}\log\mathcal{D}(\mathbf{S}_{\mathbf{Z}}^{\top})\bigg{)}$ (4)
It is evident that when $\mathbf{A}_{\mathbf{H}}$ equals the identity matrix, ReSA and InfoNCE are entirely equivalent. In other words, ReSA guides the relationship among embeddings through soft assignments obtained via clustering, whereas InfoNCE employs the identity matrix as a hard matching target, strictly enforcing the maximization of distances between all negative pairs, which may inadvertently push samples belonging to the same category further apart during training, thereby disrupting the underlying semantic cluster structure (Wang & Liu, 2021; Huang et al., 2024).
@@ 5.1 Evaluation for Classification
Evaluation on ImageNet.
Following the ImageNet evaluation protocol commonly used by SSL methods, we pretrain ResNet-50 encoders with ReSA for varying numbers of epochs. As shown in Table 2, ReSA consistently outperforms other methods on the large-scale ImageNet dataset. Remarkably, after only 100-epoch training, ReSA surpasses the performance of SimCLR, SwAV, and SimSiam trained for 800 epochs. With 200 epochs, ReSA exceeds state-of-the-art methods such as MoCoV3, Barlow Twins, VICReg, and INTL, all trained for 800 epochs. When extended to 800 epochs, ReSA achieves a linear classification accuracy of $75.2\%$, a level that methods like SwAV and DINO only reach by employing the multi-crop (Caron et al., 2020) trick. These results underscore ReSA’s exceptional potential for training on large-scale datasets. Additionally in Table 3, we conduct preliminary evaluations of ReSA’s training capability on the Vision Transformer, using a standard ViT-S/16, which has a comparable number of parameters to ResNet-50. Unlike DINO, we do not incorporate extensive training tricks, yet ReSA still outperform DINO in both linear and k-nn classification.
5.2 Analysis on Computational Overhead
[…]
Table 4: Comparison of Computational overhead among various SSL methods. For fairness, we set the batch size to 1024 with two $224^{2}$ augmented views pertraining on ImageNet, and perform all measurements including peak memory (GB per GPU) and training time (hours per epoch) on the same environment and machine equipped with 8 A100-PCIE-40GB GPUs using 32 dataloading workers under mixed-precision.

### v2

@@ Abstract
Self-supervised learning (SSL) methods via joint embedding architectures have proven remarkably effective at capturing semantically rich representations with strong clustering properties, magically in the absence of label supervision. Despite this, few of them have explored leveraging these untapped properties to improve themselves. In this paper, we provide an evidence through various metrics that the encoder’s output encoding exhibits superior and more stable clustering properties compared to other components. Building on this insight, we propose a novel positive-feedback SSL method, termed Representation Self-Assignment (ReSA), which leverages the model’s clustering properties to promote learning in a self-guided manner. Extensive experiments on standard SSL benchmarks reveal that models pretrained with ReSA outperform other state-of-the-art SSL methods by a significant margin. Finally, we analyze how ReSA facilitates better clustering properties, demonstrating that it effectively enhances clustering performance at both fine-grained and coarse-grained levels, shaping representations that are inherently more structured and semantically meaningful.
@@ 1 Introduction
Self-supervised learning (SSL) has emerged as a transformative paradigm in universal representation learning (Oord et al., 2018; Bachman et al., 2019; He et al., 2020; Chen et al., 2020a; Bao et al., 2021; Oquab et al., 2023; Assran et al., 2023), consistently surpassing supervised learning in downstream performance. Joint embedding architectures (JEA), in particular, aim to learn invariance of the same data under different transformations and noise (Bachman et al., 2019; He et al., 2020; Chen et al., 2020a), with demonstrated exceptional effectiveness in visual representation learning. Although such a pretext task may intuitively seem unrelated to capturing semantic relationships, extensive studies (Caron et al., 2018, 2021; Assran et al., 2022b) have demonstrated the strong correlation between its learned representations and semantic information.
@@
  • How to leverage clustering properties? We propose a novel SSL method, termed Representation Self-Assignment (ReSA), which employs an online self-clustering mechanism to leverage the model’s inherent clustering properties, thereby facilitating positive-feedback learning. Standard experiments demonstrate that ReSA surpasses existing state-of-the-art methods in both performance and training efficiency.
@@ 2.1 Self-Supervised Learning
In particular, the JEA usually consist of a shared encoder $E_{\theta_{e}}(\cdot)$ and projector $G_{\theta_{g}}(\cdot)$, commonly referred to as a Siamese Network (Chen & He, 2021). As shown in Figure 2, their outputs $\mathbf{H}=E_{\theta_{e}}(\mathbf{X})\in\mathbb{R}^{d_{e}\times m}$ and $\mathbf{Z}=G_{\theta_{g}}(\mathbf{H})\in\mathbb{R}^{d_{g}\times m}$ are referred to as encoding and embedding, respectively (where $m$ is the mini-batch size, $d_{e}$ and $d_{g}$ are their corresponding feature dimensions). Under this notation, we have $F_{\theta}(\cdot)=G_{\theta_{g}}(E_{\theta_{e}}(\cdot))$ with learnable parameters $\theta=\{\theta_{e},\theta_{g}\}$. It is worth noting that extensive work (Chen et al., 2020a; Gupta et al., 2022) has demonstrated that using the encoding as the representation for downstream tasks achieves much better performance than using embedding, so that the projector is only used during the pre-training process and discarded in inference.
@@
The main challenge with JEA is representation collapse, where both branches produce identical and constant outputs regardless of the inputs. Numerous paradigms have been proposed to avoid collapse, including contrastive learning methods (Chen et al., 2020a; He et al., 2020; Chen et al., 2020b, 2021; Liu et al., 2022) that attract different views from the same image (positive pairs) while pushing apart different images (negative pairs), and non-contrastive approaches (Grill et al., 2020; Caron et al., 2020, 2021; Weng et al., 2024) which directly align positive targets without incorporating negative pairs. Although  Ben-Shaul et al. (2023) had demonstrated that the encodings learned through SSL are highly correlated with semantic classes and exhibit strong clustering capabilities, few methods have leveraged this clustering ability to facilitate positive-feedback learning. A closely related work (Ma et al., 2023) exploits the encoding’s augmentation robustness to re-weight the positive alignment in the SSL objective functions; however, it still overlooks the rich clustering information inherent in the encodings.
2.2 The Information Distinction in JEA
The projector has become an indispensable component of JEA-based SSL. However, the theoretical dynamics of its optimization and the reasons behind its success remain an open question within the community. Some works have attempted to explain these principles. For example, Jing et al. (2022) empirically discovered that applying SSL loss either to encoding or embedding led to a significant decrease in the rank of the corresponding features. They argued that this rank reduction indicates a loss of diverse information, which, in turn, reduces generalization capability. This explanation aligns with the hypothesis in SimCLR (Chen et al., 2020a), where the additional projector acts as a buffer to prevent information degradation of the encoding caused by the invariance constraint. Additionally, Gupta et al. (2022)’s null space analysis for the projector posited that the projector might implicitly learn to select a subspace of the encoding, which is then mapped into the embedding. In this way, only a subspace of the encoding is encouraged to be style-invariant, while the other subspace can retain more useful information.
Therefore, the SSL constraint can cause the embedding to lose information, which may include not only clustering-irrelevant features such as background information, but also class-relevant information, making it difficult to determine which—encoding or embedding—exhibits better clustering performance in this context. In such cases, this paper first analyzes the differences between the two in terms of clustering properties empirically.
@@ 3 Exploring Clustering Properties of SSL
Ben-Shaul et al. (2023) find that within the encoder of JEA, the clustering ability of features improves progressively as intermediate layers get deeper. However, it remains unclear whether the projector exhibits a similar trend. To quantitatively evaluate the clustering performance of these components, we employ widely recognized metrics such as the Silhouette Coefficient (SC) (Rousseeuw, 1987) and Adjusted Rand Index (ARI) (Hubert & Arabie, 1985). In particular, a larger mean value of SC ($\text{SC}_{\text{mean}}$) indicates stronger local clustering ability in the representation, and a smaller standard deviation ($\text{SC}_{\text{std}}$) reflects better stability in local clustering  [^1] The ‘clustering ability’ refers to how well the vectors can represent the underlying structure of the data, while ‘stability’ signifies that the clustering results are more consistent across the data points, meaning fewer outliers and more stable cluster assignments.. Meanwhile, higher ARI values correspond to enhanced global clustering properties. The detailed introduction of these metrics can be found in the Appendix A.
@@
Figure 5: The framework of Representation Self-Assignment (ReSA). Here, no grad. denotes that the operation does not involve gradient propagation, norm signifies that each sample is $L_{2}$-normalized to compute cosine similarities, and sinkhorn refers to the Sinkhorn-Knopp algorithm used for clustering assignment.
@@ 4 Leverage Clustering Properties for Positive-Feedback Learning
Based on above analyses, we design a novel positive-feedback SSL method, which derives Representation Self-Assignment (ReSA) to guide the loss function among embeddings $\mathbf{Z}$ and $\mathbf{Z}^{\prime}$. See Figure 5 for the clear framework.
4.1 Online Self-Clustering
@@
We follow SwAV which uses only 3 iterations and sets the regularization parameter $\epsilon=0.05$. This algorithm does not involve gradient propagation, enabling it to be efficiently implemented on GPUs (Caron et al., 2020). After obtaining the doubly stochastic matrix $\mathbf{A}_{\mathbf{H}}$ as the assignment, it can naturally be utilized to guide relationship between the embeddings $\mathbf{Z}$ and $\mathbf{Z}^{\prime}$. Specially, we use the cross-entropy loss to promote the learning process:
$\ell_{\text{ReSA}}=-\frac{1}{2m}\bigg{(}$ $\sum_{i,j}\mathbf{A}_{\mathbf{H}}\circ\log\mathcal{D}(\mathbf{Z}^{\top}\mathbf{Z}^{\prime})+$ (2)
$\sum_{i,j}\mathbf{A}_{\mathbf{H}}^{\top}\circ\log\mathcal{D}(\mathbf{Z}^{\prime\top}\mathbf{Z})\bigg{)}$
where $\mathcal{D}(\mathbf{Z}^{\top}\mathbf{Z}^{\prime})=\frac{\exp(\mathbf{Z}^{\top}\mathbf{Z}^{\prime}/\tau)}{\exp(\mathbf{Z}^{\top}\mathbf{Z}^{\prime}/\tau)\mathbf{1}_{m}}$ and $\mathcal{D}(\mathbf{Z}^{\prime\top}\mathbf{Z})$ are probability distributions derived through the softmax function. $\tau$ is a scalar temperature hyperparameter, $\circ$ stands for Hadamard product, and $\mathbf{1}_{m}$ is the all-ones column vector.
@@
where $\mathbf{C}\in\mathbb{R}^{d_{g}\times K}$ is the prototype matrix learned by back-propagation, and $\mathbf{Q}=\text{sinkhorn}(\mathbf{Z}^{\top}\mathbf{C})$ is the cluster assignment using Sinkhorn-Knopp algorithm. The key differences and advantages of our ReSA compared to SwAV (and DINOv2) can be summarized as follows: (1) ReSA computes clustering assignments on encoding with high-quality clustering properties, whereas SwAV performs it on the less stable embedding. (2) SwAV requires learnable prototypes, which often necessitate complex design strategies, such as freezing prototypes during the early stages of training and use a large number of prototypes $K$ to ensure stability, whereas ReSA directly extracts clustering information from the representations. (3) SwAV executes the Sinkhorn-Knopp algorithm multiple times, corresponding to the number of global augmented views. In contrast, ReSA only requires a single execution of this algorithm regardless of the number of augmented views. This highlights the efficiency of ReSA, particularly under multi-crop scenarios.
@@
$\ell_{\text{InfoNCE}}=-\frac{1}{2m}\bigg{(}\sum_{i,i}\log\mathcal{D}(\mathbf{Z}^{\top}\mathbf{Z}^{\prime})+\sum_{i,i}\log\mathcal{D}(\mathbf{Z}^{\prime\top}\mathbf{Z})\bigg{)}$ (4)
It is evident that when $\mathbf{A}_{\mathbf{H}}$ equals the identity matrix, ReSA and InfoNCE are entirely equivalent. In other words, ReSA guides the relationship among embeddings through assignments obtained via self-clustering, whereas InfoNCE employs the identity matrix as a hard matching target, strictly enforcing the maximization of distances between all negative pairs, which may inadvertently push samples belonging to the same category further apart during training, thereby disrupting the underlying semantic cluster structure (Wang & Liu, 2021; Huang et al., 2024).
4.2 A Gradient Perspective on ReSA
Many works have conducted in-depth theoretical studies on InfoNCE, such as the alignment and uniformity properties observed by Wang & Isola (2020), the hardness-aware property discovered by Wang & Liu (2021) through gradient analysis, and the probabilistic model of InfoNCE derived by Bizeul et al. (2024) from the perspective of mutual information. Since the motivation for our proposed ReSA does not have a direct connection with the evidence lower bound of mutual information, in this section, we provide an intuitive gradient analysis to further understand the optimization mechanism of ReSA.
Given the $L_{2}$-normalized embedding vectors $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{m}]$ and $\mathbf{Z}^{\prime}=[\mathbf{z}_{1}^{\prime},\dots,\mathbf{z}_{m}^{\prime}]$, the InfoNCE formula on $\mathbf{z}_{i}$ can be writen as (omitting the symmetric terms):
$\ell_{\text{InfoNCE}}(\mathbf{z}_{i})=-\log\left(\frac{\exp(s_{i,i}/\tau)}{\sum_{k=1}^{m}\exp(s_{i,k}/\tau)}\right)$ (5)
where $s_{i,j}=\mathbf{z}_{i}^{\top}\mathbf{z}_{j}^{\prime},\forall i,j$. Defining the probability $P_{i,j}=\frac{\exp(s_{i,j}/\tau)}{\sum_{k=1}^{m}\exp(s_{i,k}/\tau)}$, the gradients of InfoNCE with respect to the positive similarity $s_{i,i}$ and the negative similarity $s_{i,j}$ ($i\neq j$) are formulated as (Wang & Liu, 2021):
$\frac{\partial\ell_{\text{InfoNCE}}(\mathbf{z}_{i})}{\partial s_{i,i}}=-\frac{1}{\tau}\sum_{k\neq i}P_{i,k},\quad\frac{\partial\ell_{\text{InfoNCE}}(\mathbf{z}_{i})}{\partial s_{i,j}}=\frac{1}{\tau}P_{i,j}$ (6)
Similarly, our ReSA formula on $\mathbf{z}_{i}$ can be writen as:
$\ell_{\text{ReSA}}(\mathbf{z}_{i})=-\sum_{j=1}^{m}{\mathbf{A}_{\mathbf{H}}}^{(i,j)}\log\left(\frac{\exp(s_{i,j}/\tau)}{\sum_{k=1}^{m}\exp(s_{i,k}/\tau)}\right)$ (7)
By contrast, the gradient of ReSA with respect to the similarity $s_{i,j}$ for any pair of samples ($\forall i,j$) takes exactly the same analytical form:
$\frac{\partial\ell_{\text{ReSA}}(\mathbf{z}_{i})}{\partial s_{i,j}}=\frac{1}{\tau}(P_{i,j}-{\mathbf{A}_{\mathbf{H}}}^{(i,j)})$ (8)
Based on the gradient analysis above, we know that InfoNCE explicitly distinguishes between the gradient forms of positive and negative similarities. This restrictive mechanism naturally leads to harmful gradient updates for negative sample pairs within the same class. In contrast, ReSA eliminates the distinction between positive and negative samples and adapts to optimize the similarity of all sample pairs by leveraging self-clustering of the encodings, thereby addressing a key challenge in contrastive learning.
@@ 5.1 Evaluation for Classification
Table 4: Comparison of Computational overhead among various SSL methods. For fairness, we set the batch size to 1024 with two $224^{2}$ augmented views pretraining on ImageNet, and perform all measurements including peak memory (GB per GPU) and training time (hours per epoch) on the same environment and machine equipped with 8 A100-PCIE-40GB GPUs using 32 dataloading workers under mixed-precision.
@@
Evaluation on ImageNet.
Following the ImageNet evaluation protocol commonly used by SSL methods, we pretrain ResNet-50 encoders with ReSA for varying numbers of epochs. As shown in Table 2, ReSA consistently outperforms other methods on the large-scale ImageNet dataset. Remarkably, after only 100-epoch training, ReSA surpasses the performance of SimCLR, SwAV, and SimSiam trained for 800 epochs. With 200 epochs, ReSA exceeds state-of-the-art methods such as MoCoV3, Barlow Twins, VICReg, and INTL, all trained for 800 epochs. When extended to 800 epochs, ReSA achieves a linear classification accuracy of $75.2\%$, a level that methods like SwAV and DINO only reach by employing the multi-crop (Caron et al., 2020) trick. These results underscore ReSA’s exceptional potential for training on large-scale datasets. Additionally in Table 3, we conduct preliminary evaluations of ReSA’s training capability on the Vision Transformer, using a standard ViT-S/16, which has a comparable number of parameters to ResNet-50. We do not incorporate extensive training tricks, yet ReSA still outperform DINO in both linear and k-nn classification.
5.2 Analysis on Computational Overhead
[…]